OpenStudy (mic123):

A candle is 17 in. tall after burning for 3 hours. After 5 hours it is 15 in. tall write a linear equation to model the relationship between height h of the candle and time T. predict how tall the candle will be after burning 8 hours.

rebeccaxhawaii (rebeccaxhawaii):

any ideas?

OpenStudy (anonymous):

Let x and y be the candles height and burn rate in inches per hour, respectively. Solve the following equations for x and y. { x - 3 y = 17, x - 5 y = 15 } // Solve x = 20, y = 1 20 - 8*1 = 12

OpenStudy (mic123):

thank you

OpenStudy (anonymous):

Your welcome and thank you for the medal.

OpenStudy (mic123):

anytime